Two rods, one made of brass and the other made of copper, are joined end to end. The length of the brass section is 0.250m and the length of the copper section is 0.750m . Each segment has cross-sectional area 6.30
Answer #1

for Copper rate of heta transfer is

PCu=kCuA(Tx-TCu)/LCu, where Tx is temperature at intersection

and for Brass

PB=kBA(TB-Tx)/LB

PCu=PB (condition for steady state transfer)

ie.kCuA(Tx-TCu)/LCu=kBA(TB-Tx)/LB

solving th ebaove equation we get

Tx=(kCuTCuLB+kBTBLCu)/(kCuLB+kBLCu)

putting all values

Tx=(400x273x0.25+109x373x0.75)/(400x0.25+109x0.75)=57792.75/181.75=317.98~318K=45oC

P=A(TB-TCu)/)(LB/KB+LCu/KCu)

P=6.3x10-3(373-273)/(0.25/109+1.75/400)=150J/Sec

since P=Q/t

Q=Pt=150x5.3x60=47700J (heat transfer in 5.3min)

also

Q=mL (where L for ice=334J/g)

m=Q/L=47700/334=142.8g of ice is melted into water by conduction of composite system